# Understanding the Reasoning Ability of Language Models From the Perspective of Reasoning Paths Aggregation

### v2

@@ Abstract
Pre-trained language models (LMs) are able to perform complex reasoning without explicit fine-tuning. To understand how pre-training with a next-token prediction objective contributes to the emergence of such reasoning capability, we propose that we can view an LM as deriving new conclusions by aggregating indirect reasoning paths seen at pre-training time. We found this perspective effective in two important cases of reasoning: logic reasoning with knowledge graphs (KGs) and math reasoning with math word problems (MWPs). More specifically, we formalize the reasoning paths as random walk paths on the knowledge/reasoning graphs. Analyses of learned LM distributions suggest that a weighted sum of relevant random walk path probabilities is a reasonable way to explain how LMs reason. Experiments and analysis on multiple KG and MWP datasets reveal the effect of training on random walk paths and suggest that augmenting unlabeled random walk reasoning paths can improve real-world multi-step reasoning performance.  [^1] We open source the code at https://github.com/WANGXinyiLinda/LM_random_walk.
@@ 2.2 Language Model Pre-training
$\small\mathcal{L}_{LM}(\theta)=\sum_{\mathcal{D}}\sum_{t=1}^{T}\log\frac{\exp{(f_{\theta}(w_{t+1}|w_{1:t}))}}{\sum_{w\in\mathcal{V}}\exp{(f_{\theta}(w|w_{1:t}))}}$ (1)
@@
$\small P_{\text{LM}}(e_{2}|e_{1},r)=\frac{\exp{(f_{\theta}(e_{2}|e_{1},r))}}{\sum_{e\in\mathcal{E}}\exp{(f_{\theta}(e|e_{1},r))}}$ (2)
@@ 2.3 Random Walk Paths Aggregation
$\small(e_{0},r,e_{n})\leftarrow(e_{0},r_{1},e_{1})\wedge(e_{1},r_{2},e_{2})\wedge...\wedge(e_{n-1},r_{n},e_{n})$
@@
$\small\mathcal{L}_{r}(w)=-\sum_{i}\left[y_{i}\ln{p_{i}}+(1-y_{i})\ln{(1-p_{i})}\right]+\lambda|w|,$
@@
$\small P_{s}(e_{2}|e_{1},r)=\frac{\exp(\sum_{h\in\mathcal{H}_{r}}P(e_{2}|e_{1},h)/T)}{\sum_{e\in\mathcal{E}}\exp(\sum_{h\in\mathcal{H}_{r}}P(e|e_{1},h)/T)}$ (4)
@@ 2.4 KL Divergence and Prediction Accuracy
$\small P_{\text{LM}}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{\text{LM}}(h|e_{1},r)$ (5)
@@
$\small P_{w}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{w}(h|e_{1},r)$ (6)
Then by the Log sum inequality, we can see that the KL divergence of the rule importance is an upper bound of the computed KL divergence  [^4] Proof available in Appendix.:
@@
[…]
@@ 2.5 Results and Analysis
We consider five KG datasets in total: Countries (Bouchard et al., 2015), UMLS (Kok & Domingos, 2007), Kinship (Denham, 2020), NELL-995 (Xiong et al., 2017), and FB15K-237 (Toutanova et al., 2015)  [^5] Dataset details can be found in Appendix.. We take the smallest two for KL divergence analysis for their lower time complexity. We show LM prediction accuracy for all datasets with different pre-training path lengths.
@@
3 Math Reasoning
After carefully analyzing the logical reasoning on KGs, we want to apply and verify the obtained insights on a more interesting and realistic case of reasoning: math reasoning with textual descriptions, which is usually called math word problems (MWPs). We continue training a pre-trained LM with random walk reasoning paths and show that these unlabeled paths consistently benefit math reasoning performance on three datasets. We also observe a similar optimal random walk path length effect as in the logical reasoning case, which is associated with the intrinsic reasoning length of different datasets. These results support our reasoning path aggregation hypothesis and imply principles for constructing/augmenting pre-training data.
@@ 3.1 Problem Setting
Suppose we have a set of training data $\mathcal{D}=\{(x^{i},r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}},y^{i})\}_{i}$, where $x^{i}$ is a question described in the text that needs to be solved by math technique. $r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}}$ is a chain-of-thought (CoT) solution, where $r^{i}_{j}$ is one reasoning step. $y^{i}$ is the ground truth answer to the question. Since MWP datasets are hard to collect and usually small in size, a model is not likely to generalize to new questions by aggregating reasoning paths over this small set of CoT reasoning paths. Fine-tuning on a pre-trained LM can effectively mitigate this problem since the LM has already seen many other reasoning paths at the pre-training time, but more unlabeled reasoning paths specific to this task would likely improve the testing performance if the path aggregation hypothesis still holds for this task.
@@ 3.2 Random Walk on Latent Reasoning Graph
We assume that CoT paths $r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}}$ can be regarded as random walk paths sampled from a reasoning graph $\mathcal{G}$, where the nodes are the reasoning states at each step $r^{i}_{j}$. The reasoning state can be regarded as a belief that will be updated after each reasoning step. Denote the last hidden state pre-trained LM we are going to tune by $f_{\theta}$. To represent the reasoning state for each step $r^{i}_{j}$, we propose to use $f_{\theta}$ to cumulatively encode all the steps before $r^{i}_{j}$, and then average over the sequence dimension, to obtain a fixed dimensional vector $s^{i}_{j}$:
$\small s^{i}_{j}=\text{avg}\;f_{\theta}(x^{i},r^{i}_{1},r^{i}_{2},...,r^{i}_{j})$
@@
3.3 Experiments
Datasets. We conduct experiments on three math word problem (MWP) datasets: GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), SVAMP (Patel et al., 2021).  [^7] Dataset details can be found in Appendix.
Training We do LoRA (Hu et al., 2021) parameter efficient training in 8 bits with Llama 2 models (Touvron et al., 2023b) because of computation limits. If not specified, we default to using the 7B version.
Results. In Table 1, we demonstrate the effectiveness of our proposed method against the supervised fine-tuning (SFT) baseline. We train both our method and SFT with $N=2500$ steps in total. The first $M=500$ steps of our method are continually pre-trained on random walk data, and then we do 2000 steps of SFT on the original dataset. Our method is about 3% better in accuracy on average over three MWP datasets and two Llama 2 models.
We also do ablation studies on two critical hyperparameters of our method: the number of steps training on random walk paths $M$ and the number of clusters/nodes $K$. In Table 2, we show that the optimal number of training steps $M$ is 500 for all three datasets. Since the generated random walk reasoning paths are not natural within small corpora, e.g. the subject might be suddenly changed from one step to another, training too many steps might make the LM overfit the unwanted artifacts. In Table 3, we show that the optimal number of clusters is 100 for all three datasets. Here 0 clusters mean the SFT baseline. Since the datasets we use are small in scale, clustering with a large number of clusters may introduce more noise than useful matchings. We hypothesize that this may be solved by using a larger dataset and more number of clusters/nodes $K$: in this case, the steps within each node will be more intrinsically similar. This also hints at the potential of our method in the actual pre-training stage: we can view each example in the pre-training corpus as a reasoning path and apply our method.
Model Method GSM8K AQUA SVAMP Avg.
7B SFT 26.8 30.0 53.3 36.7
Ours 28.5 34.6 55.8 39.6
13B SFT 37.1 35.0 66.4 46.2
Ours 41.2 37.4 69.0 49.2
Table 1: Testing accuracy of different size Llama 2 models continue pre-trained with our random walk paths and then supervised fine-tuned. The supervised fine-tuning baseline (SFT) is fine-tuned by the same number of total steps. Results are reported on three math word problem (MWP) datasets.
@@ 3.3 Experiments
#Steps GSM8K AQUA SVAMP Avg.
0 26.8 30.0 53.3 36.7
@@
Table 2: Ablation on the number of random walk training steps $M$.
#Nodes GSM8K AQUA SVAMP Avg.
0 26.8 30.0 53.3 36.7
@@
Table 3: Ablation on the number of clusters/nodes $K$.
@@ 4 Related Work
Many recent works are investigating LM’s reasoning ability. Geiger et al. (2021); Wu et al. (2023) ai to find the causal abstraction of an LM. (Hanna et al., 2023) tries to find circuit for year-span-prediction. Liu et al. (2023); Chi et al. (2023); Feng et al. (2023) show that CoTs enable fixed-size Transformers to perform certain types of reasoning tasks. Li et al. (2023); Razeghi et al. (2022); Wang et al. (2023) try to understand inference time in-context CoT reasoning. Our study is more relevant to the line of work analyzing the contribution of pre-training data to LM reasoning. Bi et al. (2023) analyzes how code data affect program-of-thoughts (Chen et al., 2023) reasoning ability. Xiao & Liu (2023); Zhou et al. (2023) study how reasoning length generalizes from training data. Ramesh et al. (2023) studies LMs’ compositional generalization ability. Our hypothesis also echos the conclusion of Malach (2023) that reasoning paths in training data enable supervision on intermediate steps with a next-token-prediction objective. Prystawski et al. (2023) propose a different hypothesis that localized structure on dependencies between variables in training data is important for LM reasoning, especially CoT reasoning. Our proposed hypothesis echoes theirs and is shown effective on more realistic data and tasks. [^8] More related work on logical reasoning and math reasoning can be found in the Appendix.
@@ Appendix A Proof
[…]
@@
$\small P_{\text{LM}}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{\text{LM}}(h|e_{1},r)$
@@ Appendix B Detailed discussion of related work
Math reasoning Recently, LM-based math reasoning models have shown to be highly effective (Azerbayev et al., 2023; Yang et al., 2023). Many works have focused on generating high-quality CoT training to improve LM’s math reasoning performance (Wang et al., 2022; Nye et al., 2022; Yuan et al., 2023). However, they all rely on the annotated Q-A pairs to generate corresponding paths with LM, which limits the size of augmented data and requires large LMs to do the CoT generation. Our proposed method does not need supervised seed data and thus can be extended to the vast amount of pre-training data. Our method is also lightweight, which only requires a small/medium LM to produce the step embeddings and then do clustering on them.
@@ C.2 Training details
Math reasoning We continually train Llama 2 (Touvron et al., 2023b) (7B and 13B parameter versions) with batch size 16 and learning rate 2e-4 using AdamW optimizer (Loshchilov & Hutter, 2017) on one 40G A100 GPU.

### v3

@@ Abstract
Pre-trained language models (LMs) are able to perform complex reasoning without explicit fine-tuning. To understand how pre-training with a next-token prediction objective contributes to the emergence of such reasoning capability, we propose that we can view an LM as deriving new conclusions by aggregating indirect reasoning paths seen at pre-training time. We found this perspective effective in two important cases of reasoning: logic reasoning with knowledge graphs (KGs) and chain-of-thought (CoT) reasoning. More specifically, we formalize the reasoning paths as random walk paths on the knowledge/reasoning graphs. Analyses of learned LM distributions suggest that a weighted sum of relevant random walk path probabilities is a reasonable way to explain how LMs reason. Experiments and analysis on multiple KG and CoT datasets reveal the effect of training on random walk paths and suggest that augmenting unlabeled random walk reasoning paths can improve real-world multi-step reasoning performance.  [^1] We open source our code at https://github.com/WANGXinyiLinda/LM_random_walk.
@@ 2.2 Language Model Pre-training
$\mathcal{L}_{LM}(\theta)=\sum_{\mathcal{D}}\sum_{t=1}^{T}\log\frac{\exp{(f_{\theta}(w_{t+1}|w_{1:t}))}}{\sum_{w\in\mathcal{V}}\exp{(f_{\theta}(w|w_{1:t}))}}$ (1)
@@
$P_{\text{LM}}(e_{2}|e_{1},r)=\frac{\exp{(f_{\theta}(e_{2}|e_{1},r))}}{\sum_{e\in\mathcal{E}}\exp{(f_{\theta}(e|e_{1},r))}}$ (2)
@@ 2.3 Random Walk Paths Aggregation
$(e_{0},r,e_{n})\leftarrow(e_{0},r_{1},e_{1})\wedge...\wedge(e_{n-1},r_{n},e_{n})$
@@
$\mathcal{L}_{r}(w)=-\sum_{i}\left[y_{i}\ln{p_{i}}+(1-y_{i})\ln{(1-p_{i})}\right]+\lambda|w|,$
@@
$P_{s}(e_{2}|e_{1},r)=\frac{\exp(\sum_{h\in\mathcal{H}_{r}}P(e_{2}|e_{1},h)/T)}{\sum_{e\in\mathcal{E}}\exp(\sum_{h\in\mathcal{H}_{r}}P(e|e_{1},h)/T)}$ (4)
@@ 2.4 KL Divergence and Prediction Accuracy
$P_{\text{LM}}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{\text{LM}}(h|e_{1},r)$ (5)
@@
$P_{w}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{w}(h|e_{1},r)$ (6)
Then by the Log sum inequality, we can see that the KL divergence of the rule importance is an upper bound of the computed KL divergence  [^4] Proof available in Appendix A.:
@@
$\text{KL}[P_{w}(\mathbf{e}|e_{1},r),P_{\text{LM}}(\mathbf{e}|e_{1},r)]$
$\leq$ $\text{KL}[P_{w}(\mathbf{h}|e_{1},r),P_{\text{LM}}(\mathbf{h}|e_{1},r)]$
@@ 2.5 Results and Analysis
We consider five KG datasets in total: Countries (Bouchard et al., 2015), UMLS (Kok & Domingos, 2007), Kinship (Denham, 2020), NELL-995 (Xiong et al., 2017), and FB15K-237 (Toutanova et al., 2015)  [^5] More dataset details can be found in Section C.1.. We take the smallest two for KL divergence analysis for their lower time complexity. We show LM prediction accuracy for all datasets with different pre-training path lengths.
@@
3 Chain-of-thoughts Reasoning
After carefully analyzing the logical reasoning on KGs, we want to apply and verify the obtained insights on a more general and realistic case of reasoning: chain-of-thoughts (CoT) reasoning (Wei et al., 2022b) with textual descriptions and step-by-step solutions. We continue training a pre-trained LM with random walk reasoning paths and show that these unlabeled paths consistently benefit CoT reasoning performance across multiple datasets of various tasks, including math reasoning, multihop question answering (QA), and logical deduction. We also observe a similar optimal random walk path length effect as in the KG logical reasoning case, which is associated with the intrinsic reasoning length of different datasets. These results support our reasoning path aggregation hypothesis and imply principles for constructing/augmenting pre-training data.
@@ 3.1 Problem Setting
Suppose we have a set of training data $\mathcal{D}=\{(x^{i},r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}},y^{i})\}_{i}$, where $x^{i}$ is a question described in the text that needs to be answered. $r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}}$ is a chain-of-thought (CoT) solution, where $r^{i}_{j}$ is one reasoning step. $y^{i}$ is the ground truth answer to the question. Since CoT datasets are hard to collect and usually small in size, a model is not likely to generalize to new questions by aggregating reasoning paths over this small set of CoT reasoning paths. Fine-tuning on a pre-trained LM can effectively mitigate this problem since the LM has already seen many other reasoning paths at the pre-training time, but more unlabeled reasoning paths specific to this task would likely improve the testing performance if the path aggregation hypothesis still holds for this task.
@@ 3.2 Random Walk on Latent Reasoning Graph
We assume that CoT paths $r^{i}_{1},r^{i}_{2},...,r^{i}_{n^{i}}$ can be regarded as random walk paths sampled from a reasoning graph $\mathcal{G}$, where the nodes are the reasoning states at each step $r^{i}_{j}$. The reasoning state can be regarded as a belief that will be updated after each reasoning step. Denote the last hidden state of the pre-trained LM we are going to tune by $f_{\theta}$. To represent the reasoning state for each step $r^{i}_{j}$, we propose to use $f_{\theta}$ to cumulatively encode all the steps before $r^{i}_{j}$, and then average over the sequence dimension, to obtain a fixed dimensional vector $s^{i}_{j}$:
$s^{i}_{j}=\text{avg}\;f_{\theta}(x^{i},r^{i}_{1},r^{i}_{2},...,r^{i}_{j})$
@@
Model Method GSM8K AQUA SVAMP StrategyQA LogicalDeduction Avg.
Gemma (2B) SFT 24.8 31.4 56.4 54.2 50.7 43.5
Ours 26.1 33.9 60.3 56.3 51.6 45.6
Yi (6B) SFT 32.2 37.0 65.8 65.8 62.2 52.6
Ours 33.1 39.8 67.0 70.0 63.3 54.6
Llama 2 (7B) SFT 26.8 30.0 53.3 58.4 55.3 44.8
Ours 28.5 34.6 55.8 63.7 56.1 47.7
Llama 2 (13B) SFT 37.1 35.0 66.4 69.5 55.7 52.7
Ours 41.2 37.4 69.0 71.2 57.7 55.3
Table 1: Testing accuracy of different open source LMs continue pre-trained with our random walk paths and then supervised fine-tuned. The supervised fine-tuning baseline (SFT) is fine-tuned by the same number of total steps. Results are reported on five CoT datasets.
@@
Ablation GSM8K AQUA SVAMP Avg.
#Steps = 0 26.8 30.0 53.3 36.7
@@
#Nodes = 0 26.8 30.0 53.3 36.7
@@
Table 2: Ablation on the number of random walk training steps $M$ and the number of clusters/nodes $K$.
3.3 Experiments
Datasets. We conduct experiments on three math word problem (MWP) datasets: GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), SVAMP (Patel et al., 2021), a multihop QA dataset StrategyQA (Geva et al., 2021), and a logical deduction dataset LogicalDeduction from the BIG-bench (Srivastava et al., 2023). GSM8K, AQUA, and SVAMP are math questions with annotated CoT steps. StrategyQA is annotated with decomposed questions, which we used as the Chain-of-thought (CoT) path of the question. As there is no CoT annotation in LogicalDeduction, we use GPT4 to generate CoTs for the training set, which on average requires 6+ reasoning steps per question.  [^7] More dataset details can be found in Section C.1.
Training Because of computation limits, we do LoRA (Hu et al., 2021) parameter efficient training in 8 bits with Llama 2 7B and 13B models (Touvron et al., 2023b), Yi 6B model (Young et al., 2024) and Gemma 2B model (Team et al., 2024). If not specified, we default to using the Llama 2 7B model.
Results. In Table 1, we demonstrate the effectiveness of our proposed method against the supervised fine-tuning (SFT) baseline. We train both our method and SFT with $N=2500$ steps in total. The first $M=500$ steps of our method are continually pre-trained on random walk data, and then we do 2000 steps of SFT on the original dataset. Experiment results show that our method can notably improve on math, multihop QA, and logical reasoning. The improvement is especially significant on StrategyQA, likely because of the relative simplicity of the reasoning, as only 3 subquestions per example on average are needed.
We also do ablation studies on two critical hyperparameters of our method: the number of steps training on random walk paths $M$ and the number of clusters/nodes $K$. In the upper half of Table 2, we show that the optimal number of training steps $M$ is 500 for all three datasets. Since the generated random walk reasoning paths are not natural within small corpora, e.g. the subject might be suddenly changed from one step to another, training too many steps might make the LM overfit the unwanted artifacts. In the lower half of Table 2, we show that the optimal number of clusters is 100 for all three datasets. Here 0 clusters mean the SFT baseline. Since the datasets we use are small in scale, clustering with a large number of clusters may introduce more noise than useful matchings. We hypothesize that this may be solved by using a larger dataset and more number of clusters/nodes $K$: in this case, the steps within each node will be more intrinsically similar. This also hints at the potential of our method in the actual pre-training stage: we can view each example in the pre-training corpus as a reasoning path and apply our method.
Figure 6: High-frequency node patterns in the training data of GSM8K and StrategyQA, discovered by our constructed latent reasoning graphs, with example CoT solutions belonging to the node pattern.
Latent reasoning graph analysis. To give a better understanding of the discovered latent reasoning graph, we show some discovered reasoning patterns through the graph in Figure 6. We show high-frequency node patterns of CoTs in the training set and corresponding CoT examples. We show examples from GSM8K and StrategyQA as they are shorter. With the GSM8K examples, we show our method discovers a 2-step pattern that first computes the baseline quantity and then performs division/multiplication to get the goal quantity based on the question specification. With the StrategyQA examples, we show a 3-step pattern that first decomposes the question into two parallel subquestions of size/weight/subject area, and then uses the third question to compare the answers to the first two questions.  [^8] More latent reasoning graph examples can be found in Section C.3.
@@ 4 Related Work
Many recent works have investigated LM’s reasoning ability. Geiger et al. (2021); Wu et al. (2023) ai to find the causal abstraction of an LM. (Hanna et al., 2023) tries to find circuit for year-span-prediction. Liu et al. (2023); Chi et al. (2023); Feng et al. (2023) show that CoTs enable fixed-size Transformers to perform certain types of reasoning tasks. Li et al. (2023); Razeghi et al. (2022); Wang et al. (2023) try to understand inference time in-context CoT reasoning. Our study is more relevant to the line of work analyzing the contribution of pre-training data to LM reasoning. Bi et al. (2023) analyzes how code data affect program-of-thoughts (Chen et al., 2023) reasoning ability. Xiao & Liu (2023); Zhou et al. (2023) study how reasoning length generalizes from training data. Ramesh et al. (2023) studies LMs’ compositional generalization ability. Our hypothesis also echos the conclusion of Malach (2023) that reasoning paths in training data enable supervision on intermediate steps with a next-token-prediction objective. Prystawski et al. (2023) propose a different hypothesis that localized structure on dependencies between variables in training data is important for LM reasoning, especially CoT reasoning. Our proposed hypothesis echoes theirs and is shown to be effective on more realistic data and tasks. Hou et al. (2023) confirm with attention probing that LMs perform multi-step reasoning internally, which echos our KG logical reasoning results.  [^9] More related work on logical reasoning and math reasoning can be found in Appendix B.
@@ 5 Conclusion
Acknowledgement
This work was supported by the National Science Foundation award #2048122. The views expressed are those of the author and do not reflect the official policy or position of the US government.
@@ Appendix A Proof
$\text{KL}[P_{w}(\mathbf{e}|e_{1},r),P_{\text{LM}}(\mathbf{e}|e_{1},r)]\leq\text{KL}[P_{w}(\mathbf{h}|r),P_{\text{LM}}(\mathbf{h}|e_{1},r)]$
@@
$P_{\text{LM}}(e_{2}|e_{1},r)=\sum_{h\in\mathcal{H}}P(e_{2}|e_{1},h)P_{\text{LM}}(h|e_{1},r)$
@@ Appendix B Detailed discussion of related work
Chain-ot-thought (CoT) reasoning Recently, LLMs have shown to be highly effective in complex reasoning tasks, like math reasoning (Azerbayev et al., 2023; Yang et al., 2023). Chain-of-thought (CoT) (Wei et al., 2022b) prompting/fine-tuning has been the major way to invoke/improve LLMs’ reasoning capabilities. Many variants of CoT prompting have been proposed to improve upon the vanilla CoT prompting (Chen et al., 2023; Yao et al., 2024). On the other hand, many works have focused on fine-tuning LLMs on generated high-quality CoT training data (Wang et al., 2022; Nye et al., 2022; Yuan et al., 2023). However, they all rely on the annotated Q-A pairs to generate corresponding paths with LM, which limits the size of augmented data and requires large LMs to do the CoT generation. Our proposed method does not need supervised seed data and thus can be extended to the vast amount of pre-training data. Our method is also lightweight, which only requires a small/medium LM to produce the step embeddings and then do clustering on them.
@@ C.1 Datasets
StrategyQA is annotated with decomposed questions, which we used as the Chain-of-thought (CoT) path of the question. Since the test set labels are not publicly released and the testing set predictions are only allowed to be verified every 7 days, we split the original training set into a new training and testing set.
@@ C.2 Training Details
CoT reasoning We continually (LORA) train all base LLMs with batch size 16 and learning rate 2e-4 using AdamW optimizer (Loshchilov & Hutter, 2017) on one 40G A100 GPU.
C.3 Additional Latent Reasoning Graph Examples
With the GSM8K examples in Figure 7, we show that our method discovers a pattern that first computes money for parallel items/individuals and then sums them up, within 3 and 4 steps respectively. With the StrategyQA example in Figure 7, we show a 2-step pattern that first asks about an emotion/psychology fact and then asks the applicability to an individual in the second question.
Figure 7: Additional high-frequency node patterns in the training data of GSM8K and StrategyQA, discovered by our constructed latent reasoning graphs, with example CoT solutions belonging to the node pattern.
Appendix D Limitations
While the scope of this project is to provide a plausible understanding of how language models obtain reasoning abilities from next-token pre-training, we acknowledge that there are other possible ways of understanding this phenomenon. While our empirical results show our hypothesis is also effective in real-world reasoning tasks, our experiments remain on a small scale with specific tasks, limited by our computation resources and project scope. An important future work is to apply our proposed random walk training method to a large and diverse reasoning corpus with more training steps in the actual pre-training phase and verify the effectiveness of our method in improving the general reasoning ability of LLMs. We also want to point out that our proposed method is effectively up-sampling the given training set and might amplify unwanted artifacts/biases if exist in the original dataset.